OpenStudy (anonymous):

t/3 - 1/2 = 1 Solve the equation. (They are fractions) t/3 - 1/2 = 1 Solve the equation. (They are fractions) @Mathematics

OpenStudy (hoblos):

t= 9/2

OpenStudy (anonymous):

\[\frac{t}{3}=1+\frac{1}{2}\]\[\frac{t}{3}=\frac{3}{2}\]\[t=\frac{9}{2}\]

OpenStudy (anonymous):

Thanks:)